# Characterizing the Sensitivity to Individual Bit Flips in Client-Side Operations of the CKKS Scheme

### v1

@@ Abstract
However, the impact of soft errors (such as bit flips) on modern HE schemes remains unexplored. Specifically, the CKKS scheme–one of the most widely used HE schemes for approximate arithmetic–lacks a systematic study of how such errors propagate across its pipeline, particularly under optimizations like the Residue Number System (RNS) and Number Theoretic Transform (NTT).
This work bridges that gap by presenting a theoretical and empirical analysis of CKKS’s fault tolerance under single bit-flip errors. We focus on client-side operations (encoding, encryption, decryption, and decoding) and demonstrate that while the vanilla CKKS scheme exhibits some resilience, performance optimizations (RNS/NTT) introduce significant fragility, amplifying error sensitivity.
By characterizing these failure modes, we lay the groundwork for error-resilient HE designs, ensuring both performance and integrity in privacy-critical applications.
@@ I Introduction
The need to preserve data privacy becomes critical when processing occurs outside the user’s trusted environment, such as in public clouds or on third-party platforms. In these scenarios, protecting data not only at rest or in transit, but also during processing, is a key challenge.
Homomorphic Encryption (HE) emerges as a promising technique to address this issue. It allows computations to be performed directly on encrypted data without needing to decrypt it first, ensuring confidentiality at all times.
@@
Unlike traditional encryption, which only protects information during storage or transmission, HE enables the secure processing of sensitive data in untrusted environments. For example, it allows a cloud server to calculate statistics on encrypted medical records or train a machine learning model using financial data without revealing its content. In these cases, the server performs the required operations without accessing the plaintext data, preserving its confidentiality throughout the process.
@@
This ability to operate directly on encrypted data is not merely convenient; it is a fundamental requirement for enabling secure computation in infrastructures where full trust cannot be assumed, such as public clouds, third-party platforms, or shared devices. Without such schemes, sensitive applications requiring delegated processing would simply be unfeasible from a privacy standpoint.
Among the most widely used homomorphic schemes is CKKS (Cheon-Kim-Kim-Song), a technique designed for operations on floating-point numbers. This makes it ideal for approximate tasks, such as numerical analysis, signal processing, or machine learning.
@@
To be considered homomorphic, an encryption scheme must be closed under at least addition and multiplication operations on encrypted data. These operations are sufficient to construct any arithmetic circuit, thereby allowing arbitrary computation without needing to decrypt the information. In addition to these fundamental operations, some modern schemes incorporate additional primitives–such as slot rotation or complex conjugation–which, while not increasing the scheme’s expressiveness, are crucial for improving efficiency and facilitating certain computational structures, especially in applications like machine learning.
HE schemes can be divided into three main categories:
@@
  • Partially Homomorphic Encryption (PHE) allows for a single type of operation on encrypted data, typically either additions or multiplications. An example is the traditional RSA[1], which supports the multiplication of encrypted data but not addition.
@@
  • Leveled Homomorphic Encryption (LHE) supports both operations (addition and multiplication), but only for a finite number of steps before the accumulated noise makes decryption returns corrupted data.
@@
  • Fully Homomorphic Encryption (FHE) enables unlimited additions and multiplications, allowing for any arbitrary computation on encrypted data.
Since Craig Gentry’s pioneering proposal [2], numerous studies have optimized the efficiency of HE schemes [3, 4, 5], improving both their software implementations and acceleration through specialized hardware [6, 7, 8].
Due to recent advancements in algorithms and optimizations, it has become feasible to ensure that a service provider cannot interpret, transfer, or access users’ encrypted data, even in the event of a security breach. In cryptographic terms, this is achieved because, with appropriate parameters, breaking a Homomorphic Encryption (HE) scheme becomes a computationally difficult problem. This reinforces confidence in the security of HE and makes its use possible in untrusted environments, such as the public cloud.
Figure 1: FHE used in a typical business application: users send their data encrypted with some homomorphic scheme to the cloud, the server operates using the scheme’s operations and returns a result that it is still encrypted.
@@
As illustrated in Fig. 1, the typical process for an FHE scheme begins with the user encrypting their data locally using their own key before sending it to the server. The server, in turn, can process this data homomorphically; that is, it performs computations directly on the encrypted data. Once processing is complete, the result, which remains encrypted, is returned to the client, who is solely responsible for its decryption.
@@
The CKKS scheme is gaining an increasing interest in machine learning and data analysis applications due to its ability to perform approximate calculations on floating-point numbers. This characteristic is fundamental for efficiently handling large data volumes and training predictive models, allowing sensitive data to be processed securely without needing decryption. Despite its high computational cost, its potential in these fields continues to be an active area of research and development [9, 10, 11].
@@
One of the most significant enhancements to the CKKS scheme is using the Residue Number System (RNS) alongside the Number Theoretic Transform (NTT). This combination has helped establish it as a standard and leads to what is known as the full RNS variant [3]. We will refer to this variant simply as CKKS, while the original version without these optimizations will be termed vanilla CKKS.
@@
It is crucial to recognize that the CKKS scheme is vulnerable to errors caused by single bit-flips, which can compromise both computational accuracy and data integrity. Since bit-flips may occur under various conditions, assessing their impact on CKKS-based applications–particularly in unreliable hardware environments–is essential.
Figure 2 demonstrates the severe impact of a single bit-flip error in a CKKS-encrypted MNIST image (implemented using OpenFHE, a widely adopted FHE library). Despite the ciphertext consisting of millions of bits under standard parameters, flipping just one bit leads to a complete failure in image reconstruction, as evidenced by the corrupted output in Fig. 2(b).
@@
In this work, we conduct a comprehensive analysis of the CKKS scheme’s sensitivity to single-bit errors in client-side stages, such as encoding, encryption, decryption, and decoding operations. We focus on the impact of RNS (Residue Number System) and NTT (Number Theoretic Transform) optimizations on the scheme’s resilience against these errors, exploring an understudied area in the homomorphic encryption literature.
Unlike other proposals focused on specific applications, our analysis provides a general perspective aimed at establishing a clear framework to enhance the security and robustness of CKKS in diverse contexts. Our main contributions are as follows:
@@
  • We provide a generalized framework that can serve as a foundation for future research aimed at improving the security and resilience of CKKS against hardware failures.
In the rest of the paper, Section II describes in detail the CKKS scheme in its standard version and its full variant with RNS and NTT. Next, Section III presents a theoretical analysis of CKKS’s robustness against bit-flip errors. Then, Section IV details the methodology used in our experimental campaigns. The experimental analysis is developed in Section V, and finally, Section VI presents the conclusions of the work.
@@ II Background and CKKS Overview
In this section, we first introduce the core parameters of the scheme, followed by a conceptual overview alongside a simplified explanation of its key components–making the material accessible even to readers without a specialized background in cryptography.
We then delve into the mathematical foundations required for the analysis presented in Section III. Readers primarily interested in the experimental methodology may proceed directly to Section IV.
@@ II-A General Flow of the CKKS Scheme
CKKS is an approximate homomorphic encryption scheme designed to enable computations on encrypted data while providing explicit control over approximation error. It operates over a polynomial ring of degree $N$ and is based on the Ring Learning With Errors (RLWE) problem, a well-established lattice-based problem that ensures security through the controlled injection of noise into ciphertexts.
CKKS supports the encrypted processing of vectors of real or complex numbers, without ever revealing the original data.
Figure 1 presents the FHE’s process consisting of five main stages:
@@
[…]
@@
[…]
    $\boldsymbol{z}=(P(\xi),P(\xi^{3}),\cdots,P(\xi^{2N-1})).$
@@
    Decryption and Decoding: The ciphertext is multiplied with the secret key and summed. The result is then divided by $\Delta$, and a specialized FFT is applied to recover an approximation of the original input vector.
Since the security of CKKS relies on the computational hardness of the RLWE problem, each encryption inherently introduces noise. Although it is small at the beginning, it accumulates during homomorphic operations. In general, each operation increases the noise level in the ciphertext, with homomorphic multiplication contributing the most to this growth.
For this reason, the number of successive multiplications is often used as an indirect estimate of the accumulated error in a ciphertext. If this growth is not properly managed–e.g., through techniques such as bootstrapping–the error can exceed tolerable limits, resulting in incorrect outputs during decryption and decoding.
@@ II-B Parameters controlling CKKS
where $\mathbb{Z}_{q}$ denotes the integers modulo $q$, that is, the integer values between 0 and $q-1$ (addition and multiplication are performed modulo $q$), $\mathbb{Z}_{q}[X]$ denotes the polynomials with coefficients in $\mathbb{Z}_{q}$, and $/(X^{N}+1)$ indicates that polynomials are considered modulo $X^{N}+1$.
@@
This ring imposes a negacyclic structure that enables, after encoding, the efficient representation of vectors of complex numbers. In CKKS, data is encoded as a polynomial that, when evaluated at specific values, yields each of the encoded data points. In particular, the degree N must be a power of two and serves two critical purposes: it determines the maximum number of complex values that can be encoded simultaneously ($N/2$), and it is essential for maintaining the security of the scheme.
@@
The original (i.e., vanilla) version of the CKKS scheme uses the following core parameters, which are selected to achieve the desired security level, maintain tolerable error growth, and minimize computational cost:
[…]
@@
  • $q_{0}$ is the initial modulus—a single large integer (typically up to 60 bits)–that largely defines the target precision.
@@
  • $\Delta$ is the scaling factor, a constant used to amplify real values before encoding in order to preserve floating-point precision.
@@
  • $L$ is the multiplicative depth, that is, the number of homomorphic multiplications that can be performed before accumulated noise degrades precision.
@@
The total modulus of the scheme, denoted $Q$, and the size of the coefficients in the encoding and encryption polynomials are defined differently depending on the CKKS variant used. In the vanilla version, $Q$ is given by:
$Q=\Delta^{L}\cdot q_{0},$
@@
In CKKS, each multiplication between ciphertexts increases both the error and the magnitude of the encoded values. To control this growth and preserve precision, an operation called rescaling is applied. Rescaling reduces the scale of the encoded values and also decreases the active modulus, which in turn limits the number of consecutive multiplications (with rescaling) that can be performed. This limit is known as the multiplicative depth of the scheme. The parameter $\Delta$ is applied $L$ times–once for each multiplicative level supported by the modulus hierarchy.
II-B1 CKKS full RNS variant
The Residue Number System (RNS) is an alternative representation of integers that expresses a number through its residues modulo several small, pairwise coprime integers. This encoding allows arithmetic operations–such as addition and multiplication–to be performed independently and in parallel on each modulus, reducing computational complexity and significantly improving efficiency. These properties make RNS particularly well-suited for high-performance contexts such as homomorphic encryption.
The use of the Chinese Remainder Theorem (CRT) is essential in RNS, as it ensures that each set of residues uniquely corresponds to a single integer within a defined range. Moreover, CRT provides an efficient way to reconstruct the original integer from its residue representation.
In the RNS-optimized variant of the CKKS scheme, all polynomials involved (i.e., messages, keys, and ciphertexts)are represented as multiple polynomials of the same degree, one for each residue modulo in the RNS base. Additionally, each of these polynomials is transformed using the Number Theoretic Transform (NTT)–a modular analogue of the Fast Fourier Transform (FFT). NTT accelerates polynomial multiplication, which is one of the most computationally intensive operations in the scheme.
@@ II-B1 CKKS full RNS variant
These optimizations–RNS and NTT–do not alter the semantics of the encryption scheme but do change its internal implementation. In particular, they have implications for error propagation and sensitivity, topics that will be analyzed in detail in Section III.
In this variant, the total modulus of the scheme is expressed as:
@@
where $q_{0}$ is the initial modulus and $q_{1},\dots,q_{L}$ are additional moduli that in general are selected by the scheme, all of similar bit length. It is worth noting that some libraries provide manual control over the bit-length of these $q_{i}$ moduli, the implementations employed in this work automatically configure these parameters according to the bit length specified for $q_{0}$.
II-C CKKS’ Insights
The CKKS scheme accepts as input a vector of up to $N/2$ floating-point elements, i.e., $\in\mathbb{C}^{N/2}$. This input is mapped into the polynomial ring $\mathbb{Z}[X]/(X^{N}+1)$ through a process known as encoding, producing a plaintext polynomial $m$. Here, $\mathbb{Z}[X]$ denotes the ring of polynomials with integer coefficients, and $\mathbb{Z}[X]/(X^{N}+1)$ is the quotient ring where polynomials are reduced modulo $X^{N}+1$. This implies working with polynomials of degree at most $N-1$ under negative (or negacyclic) convolution. In this context, polynomial reduction is performed via the substitution $X^{N}=-1$, ensuring that results remain within the allowed polynomial space of degree less than $N$.
The resulting plaintext is then encrypted using either the public or secret key, following the RLWE problem, where a small amount of noise is intentionally injected for security. The ring $\mathbb{Z}[X]/(X^{N}+1)$ is preferred in RLWE-based schemes due to its stronger security properties compared to $\mathbb{Z}[X]/(X^{N}-1)$, where the convolution is cyclic and considered more vulnerable to attacks.

### v3

@@ Abstract
However, the impact of soft errors (such as bit-flips) on modern HE schemes remains unexplored. Specifically, the CKKS scheme–one of the most widely used HE schemes for approximate arithmetic–lacks a systematic study of how such errors propagate across its pipeline, particularly under optimizations like the Residue Number System (RNS) and Number Theoretic Transform (NTT).
This work bridges this gap by presenting a theoretical and empirical analysis of CKKS’s fault tolerance behavior under bit-flip errors. We focus on client-side operations (encoding, encryption, decryption, and decoding) and demonstrate that while the vanilla CKKS scheme exhibits some resilience, performance optimizations like RNS and NTT exacerbate the error sensitivity of the scheme. By characterizing these failure modes, we lay the groundwork for developing HE hardware and software systems that are resilient to errors, ensuring both performance and integrity in privacy-critical applications.
@@ I Introduction
The need to preserve data privacy becomes critical when processing occurs outside the client’s trusted environment, such as in public clouds or on third-party platforms. In these scenarios, protecting data not only at rest or in transit, but also during processing, is a key challenge. Homomorphic Encryption (HE) emerges as a promising technique to address this issue by enabling computation on encrypted data (i.e., without requiring decryption), thereby ensuring confidentiality at all times. For example, it allows a cloud server to calculate statistics on encrypted medical records or train a machine learning model using financial data without revealing its content. In these cases, the server performs the required operations without accessing the plaintext data, preserving its confidentiality throughout the process.
@@
This ability to operate directly on encrypted data is not merely convenient; it is a fundamental requirement for enabling secure computation in infrastructures where full trust cannot be assumed, such as public clouds, third-party platforms, or shared devices. Without such schemes, sensitive applications requiring delegated processing would be unfeasible from a privacy standpoint. Among the most widely used homomorphic schemes is CKKS (Cheon-Kim-Kim-Song) [1] due to its fixed-point arithmetic support. This makes it ideal to handle real number computation in tasks such as numerical analysis, signal processing, or machine learning.
Figure 1: Typical HE use case: a client sends encrypted data to the cloud, the cloud operates on the encrypted data and returns the results. Data and results remain encrypted all the time. [^1] Adapted from https://openmined.org/blog/from-fully-homomorphic-encryption-to-silicon/
@@
An encryption scheme should support encrypted computation of at least addition and multiplication operations in order to be considered homomorphic. These operations are sufficient to construct any arithmetic circuit, thereby allowing arbitrary computation without decrypting the information. In addition to these fundamental operations, some modern schemes incorporate additional primitives–such as elements rotation or complex conjugation–which, while not increasing the scheme’s expressiveness, are crucial for improving efficiency and facilitating certain computational structures, especially in applications like machine learning. HE schemes can be divided into three main categories:
@@
  • Partial Homomorphic Encryption (PHE) supports a single type of operation on encrypted data, either addition or multiplication. An example is The RSA (Rivest–Shamir–Adleman) cryptosystem[2], which supports multiplication of encrypted data but not addition.
@@
  • Leveled Homomorphic Encryption (LHE) supports both operations (addition and multiplication), but only for a limited number of steps before the accumulated noise makes correct decryption infeasible.
@@
  • Fully Homomorphic Encryption (FHE) enables unlimited additions and multiplications, enabling arbitrary computation on encrypted data.
@@
As illustrated in Figure 1, the typical process for an HE scheme begins with the client encrypting data locally using a secret key before sending it to the cloud server. The server process this data homomorphically; that is, it performs computations directly on the encrypted data. Once processing is complete, the result, which remains encrypted, is returned to the client, who is solely responsible for its decryption.
@@
The CKKS scheme is gaining increasing interest in machine learning and data analysis applications due to its ability to perform approximate calculations on floating-point numbers. This characteristic is fundamental for efficiently handling large data volumes and training predictive models, allowing sensitive data to be processed securely without needing decryption. Despite its high computational cost, its potential in these fields is still an active area of research and development [3, 4, 5].
@@
One of the most significant enhancements to the CKKS scheme is using the Residue Number System (RNS) alongside the Number Theoretic Transform (NTT). This combination has helped establish it as a standard and leads to what is known as the full RNS variant [6]. Throughout this paper, we will refer to this variant simply as CKKS, while the original version without these optimizations will be termed vanilla CKKS.
@@
It is crucial to recognize that the CKKS scheme is vulnerable to errors caused by single bit-flips, which can compromise both computational accuracy and data integrity. Since bit-flips may occur under various conditions, assessing their impact on CKKS-based applications–particularly in unreliable hardware environments–is crucial. Figure 2 illustrates a scenario in which an image from the MNIST dataset (Figure 2(a)) is processed through the CKKS pipeline. Figure 2(b) shows an entirely distorted image recovered when a single bit error is introduced during the encoding phase using RNS and NTT.
@@
In this work, we conduct a comprehensive analysis of the CKKS scheme’s sensitivity to single bit errors in client-side stages, such as encoding, encryption, decryption, and decoding operations. We focus on the impact of RNS (Residue Number System) and NTT (Number Theoretic Transform) optimizations on the scheme’s resilience against these errors. To the best of our knowledge, this paper is the first study that explores this aspect of HE. Our analysis provides a general framework aimed at enhancing the security and robustness of CKKS in diverse contexts. Our main contributions are as follows:
@@
  • We provide a framework that can serve as a foundation for future research aimed at improving the security and resilience of CKKS against hardware failures.
Section II describes the CKKS scheme in its vanilla and full variants. Section III presents a theoretical analysis of CKKS’s robustness against bit-flip errors. Section IV details the methodology used in our experimental campaigns. Section V develops the experimental analysis and discusses our findings. Finally, Section VI presents our conclusions.
@@ II Background and CKKS Overview
In this section, we first introduce the core parameters of the CKKS scheme, followed by a conceptual overview alongside a simplified explanation of its key components–making the material accessible even to readers without a specialized background in cryptography. We then delve into the mathematical foundations required for the analysis presented in Section III.
@@ II-A General Flow of the CKKS Scheme
CKKS is an approximate homomorphic encryption scheme designed to enable computations on encrypted data while providing explicit control over approximation error. It operates over a polynomial ring of degree $N$ based on the Ring Learning With Errors (RLWE) problem, a lattice-based problem that ensures security through the controlled injection of noise into ciphertexts. Figure 1 shows the processing stages:
@@
    Key Generation: The secret key and public key are generated as degree-$N$ polynomials, created through random sampling and error injection.
@@
    Encoding: The input vector is transformed using a specialized inverse Fast Fourier Transform (IFFT) into a degree-$N$ polynomial, constructed such that its evaluations at certain roots of unity approximately reproduce the original values. The resulting polynomial’s coefficients are then scaled by $\Delta$ and rounded to integers, producing the plaintext. Formally, given a polynomial $P$ and a vector $z\in\mathbb{R}^{N}$, the following relation holds:
    $\boldsymbol{z}\approx(P(\xi),P(\xi^{3}),\cdots,P(\xi^{2N-1})).$
@@
    Decryption and Decoding: One component of the ciphertext is multiplied by the secret key and added to the other component. The result is a plaintext, which is divided by $\Delta$, and then a specialized FFT is applied to recover an approximation of the original input vector.
Since the security of CKKS relies on the computational hardness of the RLWE problem, each encryption inherently introduces noise. Although it is small at the beginning, it accumulates during homomorphic operations. Each operation increases the noise level in the ciphertext, with homomorphic multiplication contributing the most to this growth. For this reason, the number of successive multiplications is often used as an indirect estimate of the accumulated error in a ciphertext. If this growth is not properly managed–e.g., through techniques such as bootstrapping [14]–the error can exceed tolerable limits, resulting in incorrect outputs during decryption and decoding.
@@ II-B Parameters controlling CKKS
where $\mathbb{Z}_{q}$ denotes the integers modulo $q$ (i.e., the integer values between 0 and q1q-1, with addition and multiplication performed modulo $q$)); $\mathbb{Z}_{q}[X]$ denotes the polynomials with coefficients in $\mathbb{Z}_{q}$; and $(X^{N}+1)$ indicates that polynomials are considered modulo $X^{N}+1$.
@@
This ring imposes a negacyclic structure that enables, after encoding, the efficient representation of vectors of complex numbers. In CKKS, data is encoded as a polynomial that, when evaluated at specific values, yields each of the encoded data points. In particular, the degree $N$ must be a power of two and serves two critical purposes: it determines the maximum number of complex values that can be encoded simultaneously, and enforces the security of the scheme.
@@
The original (vanilla) version of the CKKS scheme uses the following core parameters, which are selected to achieve the desired security level, maintain tolerable error growth, and minimize computational cost:
  • NN is the ring degree, typically in the range $2^{12}-2^{16}$, and it determines both the dimension of the encoding space and the computational cost.
@@
  • $q_{0}$ is the initial modulus–a single large integer (typically up to 60 bits)–that determines the target precision.
@@
  • $\Delta$ is the scaling factor, a constant used to amplify real values before encoding to preserve floating-point precision.
@@
  • $L$ is the multiplicative depth; i.e., the number of homomorphic multiplications that can be performed before accumulated noise degrades precision.
@@
The total modulus of the scheme, denoted $Q$, and the size of the coefficients in the encoding and encryption polynomials are defined differently depending on the CKKS variant used. In the vanilla version, $Q=\Delta^{L}\cdot q_{0}$.
@@
In CKKS, each multiplication between ciphertexts increases both the error and the magnitude of the encoded values. To control this growth and preserve precision, an operation called rescaling is applied. Rescaling reduces the scale of the encoded values and also decreases the active modulus, which in turn limits the number of consecutive multiplications (with rescaling) that can be performed. This limit is known as the multiplicative depth of the scheme. The factor $\Delta$ appears once at each of the $L$ multiplicative levels of the modulus hierarchy, which explains the term $\Delta^{L}$ in the expression for $Q$.
II-C Full CKKS Variant
The magnitude of the modulus $Q$ (and consequently, the size of the polynomial coefficients) can range from hundreds to thousands of bits, depending on the use case. However, modern computing systems typically operate with word sizes of 64 bits or less, making direct operations on such large coefficients impractical. The Residue Number System (RNS) enables the representation of integer values using their remainders with respect to a set of pairwise coprime moduli. This allows arithmetic operations–such as addition and multiplication–to be performed using 64-bit word size, reducing computational complexity and significantly improving efficiency. These properties make RNS particularly well-suited for high-performance contexts such as homomorphic encryption. The use of the Chinese Remainder Theorem (CRT) is essential in RNS, as it ensures that each set of residues uniquely corresponds to a single integer within a defined range. Moreover, CRT provides an efficient procedure to reconstruct the original integer from its residue representation.
Once mapped to the Residue Number System (RNS), each polynomial is further transformed via the Number Theoretic Transform (NTT)–the modular analogue of the Fast Fourier Transform (FFT). This reduces the complexity of polynomial multiplication (one of the most computationally demanding operations in the scheme) from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$. By default, the CKKS scheme assumes that polynomials are represented in the evaluation domain (NTT). When certain operations require polynomials to be in the coefficient domain, an inverse NTT (iNTT) is applied to transform the polynomial from the evaluation domain back to the coefficient domain.
@@ II-C Full CKKS Variant
These optimizations (RNS and NTT) do not alter the semantics of the encryption scheme but do change its internal implementation. In particular, they have implications for error propagation and sensitivity, topics that will be analyzed in detail in Section III.
In the full CKKS variant, the total modulus of the scheme is expressed as:
@@
where $q_{0}$ is the initial modulus and $q_{1},\dots,q_{L-1}$ are additional moduli that in general are selected by the scheme, all of similar bit length. It is worth noting that some libraries provide manual control over the bit-length of these $q_{i}$ moduli; the implementations employed in this work automatically configure these parameters according to the bit length specified for $q_{0}$.
II-D Encoding/Encryption/Decryption/Decoding
The CKKS scheme operates on complex vectors of size up to $N/2$. This input is encoded into the polynomial ring $\mathbb{Z}[X]/(X^{N}+1)$, producing a plaintext polynomial $m$. The resulting plaintext is then encrypted using either the public or secret key, following the RLWE problem, where a small amount of noise is intentionally injected for security. The ring $\mathbb{Z}[X]/(X^{N}+1)$ is preferred in RLWE-based schemes due to its stronger security properties compared to $\mathbb{Z}[X]/(X^{N}-1)$, where the convolution is cyclic and considered more vulnerable to attacks.